# A Digital twin for Diesel Engines: Operator-infused PINNs with Transfer Learning for Engine Health Monitoring

### v1

@@ Appendix C Additional Tables
Table C.1: Input output structure of DeepONets for EGR ($\tilde{u}_{egr1}$, $\tilde{u}_{egr1}$) and VGT ($\tilde{u}_{vgt}$) states. As shown in Fig. 5, the DeepONet ($\mathcal{D}_{1}$) for EGR states has four branches and two outputs. The DeepONet ($\mathcal{D}_{1}$) for VGT state has two branches and one output. The input to the trunk is $t=0-60$ sec and scaled between $[-1,1]$. A schematic diagram of the DeepONet is shown in Fig. 5.
@@
$\begin{matrix}t=t_{0}=0.00\\ t=t_{1}=0.02\\ \cdot\\ \cdot\\ t=t_{n}=60.0\end{matrix}$ $\begin{matrix}u_{degr,0}&0&0&\dots&0\\ u_{degr,1}&u_{degr,0}&0&\dots&0\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ u_{degr,n}&u_{degr,n-1}&u_{degr,n-2}&\dots&u_{degr,0}\\ \end{matrix}$ $\begin{matrix}\tilde{u}_{egr1,0}\\ \tilde{u}_{egr1,0}\\ \cdot\\ \cdot\\ \tilde{u}_{egr1,0}\end{matrix}$ $\begin{matrix}u_{degr,0}&0&0&\dots&0\\ u_{degr,1}&u_{degr,0}&0&\dots&0\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ u_{degr,n}&u_{degr,n-1}&u_{degr,n-2}&\dots&u_{degr,0}\\ \end{matrix}$ $\begin{matrix}\tilde{u}_{egr2,0}\\ \tilde{u}_{egr2,0}\\ \cdot\\ \cdot\\ \tilde{u}_{egr2,0}\end{matrix}$ $\begin{matrix}\tilde{u}_{egr1,0}\\ \tilde{u}_{egr1,1}\\ \cdot\\ \cdot\\ \tilde{u}_{egr1,n}\end{matrix}$ $\begin{matrix}\tilde{u}_{egr2,0}\\ \tilde{u}_{egr2,1}\\ \cdot\\ \cdot\\ \tilde{u}_{egr2,n}\end{matrix}$
@@
$\begin{matrix}t=t_{0}=0.00\\ t=t_{1}=0.02\\ \cdot\\ \cdot\\ t=t_{n}=60.0\end{matrix}$ $\begin{matrix}u_{dvgt,0}&0&0&\dots&0\\ u_{dvgt,1}&u_{dvgt,0}&0&\dots&0\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ u_{dvgt,n}&u_{dvgt,n-1}&u_{dvgt,n-2}&\dots&u_{dvgt,0}\\ \end{matrix}$ $\begin{matrix}\tilde{u}_{vgt,0}\\ \tilde{u}_{vgt,0}\\ \cdot\\ \cdot\\ \tilde{u}_{vgt,0}\end{matrix}$ NA NA $\begin{matrix}\tilde{u}_{vgt,0}\\ \tilde{u}_{vgt,1}\\ \cdot\\ \cdot\\ \tilde{u}_{vgt,n}\end{matrix}$

### v2

@@ Appendix D Detailed loss function for engine model
Table A-1: Input output structure of DeepONets for EGR ($\tilde{u}_{egr1}$, $\tilde{u}_{egr1}$) and VGT ($\tilde{u}_{vgt}$) states. As shown in Fig. A-2, the DeepONet ($\mathcal{D}_{1}$) for EGR states has four branches and two outputs. The DeepONet ($\mathcal{D}_{1}$) for VGT state has two branches and one output. The input to the trunk is t=060t=0-60 sec and scaled between [1,1][-1,1]. A schematic diagram of the DeepONet is shown in Fig. A-2.
@@
[…]
@@
[…]
Table A-2: DeepONet results for independent variables: The training and testing errors for DeepONet for EGR and VGT states. The two DeepONets are trained using clean data and with added Gaussian noise. The relative $L_{2}$ error (%) for training and testing are shown in the fourth and sixth columns, respectively. The bottom part of the table gives the $L_{2}$ error (%) in training and testing when Gaussian noise is added to the training labelled data. We also note that the error in the case of noisy data is calculated with respect to the clean data. The mean and standard deviation of the error are calculated based on six different independent runs.
DeepONet Output Training Testing
Variable Data type Rel. $L_{2}$ error (%) Data type Rel. $L_{2}$ error (%)
$\tilde{u}_{egr2}$ Clean data $0.65\pm 0.04$ Clean data $3.64\pm 0.18$
with added noise in output during training
$\tilde{u}_{egr2}$ 3% noise added $0.97\pm 0.02$ Clear data $3.48\pm 0.17$
Appendix F Additional results for Multi-stage transfer learning prediction
In this section, we provide additional results showing the predictions for different engine related Quantities of Interest (QOIs). Fig. A-3 show the variation in parameter estimation over 30 independent runs with 301 data points under clean conditions. Adding residual points help in reducing the variation in estimates in some scenarios. This, however comes at an additional computational cost. Higher variation is observed in time segment 1-2 minute for all parameters likely due the engine entering idling condition from a dynamical state which induces bias in PINN training. Figure A-4 and A-5 show prediction results for main state estimates and flow rates obtained from one instance of the model run. Additional Figures A-6 - A-11 show the ensemble plot for different predictions over the 30 independent runs of the model for both clean and noisy test conditions. We notice higher variation in predictions with noisy data for all dynamic estimates. Overall, the mean estimated values are reasonably close to the ground truth values.
Fig. A-3: Violin plot, multi-stage TL, clean data, 301 data points: variation in engine parameter predictions over 30 runs with clean data and 301 data points. More data points help reduce the variation in estimation across all data segments as compared to the model trained with 151 data points
Fig. A-4: Predicted states, multi-stage TL, clean data, 151 data points: Primary gas flow dynamics states from multi-stage transfer learning model (data scale normalized due to confidentiality). The plot shows a representative case where clean data is used with 151 data points for online training. Note that states $\tilde{u}_{egr1},\tilde{u}_{egr2},$ and $\tilde{u}_{vgt}$ are outputs obtained from pre-trained DeepONet and are not trained during the transfer learning process.
Fig. A-5: Gas flow rates, multi-stage transfer learning, clean data, 151 data points: Gas flow rates from transfer learning model. The plot shows a representative case where clean data is used with 151 data points for online training.
Fig. A-6: Predicted dynamic states, multi-stage TL, clean data, 151 data points, ensemble: Ensemble prediction results for primary gas flow dynamics states from transfer learning model (data scale normalized due to confidentiality). The plot shown is for scenario where clean data is used with 151 data points used for online training. The ensemble mean (red) for the 30 runs is plotted against ground truth (black) for the eight dynamic states defined by ODE’s describing the gas flow dynamics. Error bands (shown in light green) indicate the spread of data between point-wise maximum and minimum values across all 30 runs.
Fig. A-7: Predicted gas flow, multi-stage TL, clean data, 151 data points, ensemble: Ensemble prediction results for gas flow rates from the Transfer learning model with clean data over 30 independent runs across all segments. 151 data points were used during online parameter estimation task.
Fig. A-8: Predicted empirical variables, multi-stage TL, clean data, 151 data points, ensemble: Ensemble prediction estimates for empirical relationships for clean data with Transfer learning model. Note that in region $\approx 1.5-3.5\,min$, the model’s prediction indicates a spectral bias in learning the high frequency signals that typically result from engine idling. Elsewhere, the model predictions show good match with target values for these empirical relationships.
Fig. A-9: Predicted dynamic states, multi-stage TL, noisy data, 301 data points, ensemble: Ensemble prediction results for primary gas flow dynamics states from Transfer learning model (data scale normalized due to confidentiality). The plot shown is for scenario where noisy data is used with 301 data points used for online training. The ensemble mean (red) for the 30 runs is plotted against ground truth (black) for the eight dynamic states defined by ODE’s describing the gas flow dynamics.
Fig. A-10: Predicted gas flow rate, multi-stage TL, noisy data, 301 data points, ensemble: Ensemble prediction results for gas flow rates from the Transfer learning model with noisy data. 301 data points were used during online parameter estimation task.
Fig. A-11: Predicted empirical variable, multi-stage TL, noisy data, 301 data points, ensemble: Ensemble prediction estimates for empirical relationships for noisy data with transfer learning model with 301 data points. Note that in region $\approx 1.5-3.5\,min$, the model’s prediction indicate a spectral bias in learning the high frequency signals that typically result from engine idling. Also, there is a significant increase in expected variability in prediction of these empirical states under noisy conditions as compared to clean conditions shown in figure A-8, correlating with increase in variability in prediction of other states and parameters as well.
Appendix G Additional results for Few shot transfer learning
In this section, we provide additional results showing the predictions of different quantity for few-shot transfer learning method. In Table A-3, we have shown the training of multi-head neural network for few-shot TL. As discussed in the main text, these networks are trained offline using data and Adam [41] optimizer. One of the realizations of the state variable and $x_{r}$ and $T_{1}$ are shown in Fig. A-13 and intermediate variables are shown in Fig. A-14 for the clean data case with 151 data points. As discussed in the main text, we also consider a case with 301 clean data. The violin plots of the predicted unknowns are shown in Fig. A-12. The ensemble plots for the different variables corresponding to 151 clean data case are shown in Figs. A-15, A-16 and A-17. The corresponding ensemble plots for different variables in the case of 301 noisy data points are shown in Figs. A-18, A-19 and A-20
Table A-3: Training of multi-head networks for few-shot transfer learning (Phase-I): Multi-head neural networks considered to approximate the state variables and $x_{r}$ and $T_{1}$. The input to the network is time ($t$). The input time is 0 to 60 sec with $\delta t=0.2$ sec, and the input is scaled between [1,1][-1,1]. Each head approximates a minute of the variable. We considered 240 heads for each output approximating 240 minutes of dynamics. Thus, network $\mathcal{N}_{2}(t,\bm{\theta}_{2})$, $\mathcal{N}_{3}(t,\bm{\theta}_{3})$ and $\mathcal{N}_{4}(t,\bm{\theta}_{4})$ has 240 output, however network $\mathcal{N}_{1}(t,\bm{\theta}_{1})$ has 480 output as it approximate two ($p_{em}$ and $p_{im}$) variables. The ”Output transformation” and scaling are considered as per Table 4. The $L_{2}$ error shows % relative $L_{2}$ error in training for each variable.
Network Output Rel. $L_{2}$ error (%)
$\mathcal{N}_{1}(t,\bm{\theta}_{1})$ $p_{em}$, $p_{im}$ $p_{em}:1.478\pm 0.135$, $p_{im}:1.476\pm 0.078$
$\mathcal{N}_{2}(t,\bm{\theta}_{2})$ $x_{r}$ $5.512\pm 0.136$
$\mathcal{N}_{3}(t,\bm{\theta}_{3})$ $T_{1}$ $0.913\pm 0.008$
$\mathcal{N}_{4}(t,\bm{\theta}_{4})$ $\omega_{t}$ $1.035\pm 0.028$
Fig. A-12: Violin plot, few-shot learning, clean data, 301 data points: of the predicted value of the unknown parameters. 301 data points are considered for each of the known parameters (for each one minute segment) and no noise is considered here.
Fig. A-13: Predicted states, Few-shot learning, Clean data, 151 data points: One of the predicted realization of the state variables and $x_{r}$ and $T_{1}$, for the case where clean data and 151 data points are considered for each known variables (for each one minute duration). Each 1 minute is trained independently and combined to get the total 5 minutes results. The vertical dotted blue lines separate each minute result. The results are normalized using Eq. (21) with maximum and minimum true value calculated over the 5 minute.
Fig. A-14: Intermediate variables, Few-shot learning, Clean data, 151 data points: One of the predicted realization of different variables of the diesel engine for the case where clean data and 151 data points are considered for each known variable (for each one minute duration). The variable $A_{egr}$, $T_{e}$, $T_{em}$ and $W_{t}$ are the variable which are directly depends on the unknown parameters $A_{egrmax}$, $\eta_{sc}$, $h_{tot}$ and $A_{vgtmax}$. The variable $T_{e}$ also depends on the variable $T_{1}$ and $x_{r}$.
Fig. A-15: Predicted states, Few-shot learning, Clean data, 151 data points, ensemble: Ensemble prediction for state variable and $x_{r}$ and $T_{1}$ for clean data with 151 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Fig. A-16: Predicted states, Few-shot learning, Clean data, 151 data points, ensemble: Ensemble prediction for different variables for clean data with 151 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Fig. A-17: Predicted empirical formulae, Few-shot learning, Clean data, 151 data points, ensemble: Ensemble prediction for the empirical formulas for clean data with 151 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Fig. A-18: Predicted states, Few-shot learning, Noisy data, 301 data points, ensemble: Ensemble prediction for state variable and $x_{r}$ and $T_{1}$ for noisy data with 301 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Fig. A-19: Intermediate variables, Few-shot learning, Noisy data, 301 data points, ensemble: Ensemble prediction for different variables for noisy data with 301 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.
Fig. A-20: Predicted empirical variables, Few-shot learning, Noisy data, 301 data points, ensemble: Ensemble prediction for the empirical formulas for noisy data with 301 data points in the data loss predicted using few-shot transfer learning. The ensembles are calculated over 30 independent runs for each time segment.